# Correction to Hawking radiation in non-singular gravitational collapse

### v2

@@ II Black-hole dynamics
This concludes our review of the tools required to study dynamical black holes. Next we review a framework in which spontaneous emission by such black holes can be conveniently described.
@@ III Quantum theory of a relativistic particle
$\Box\phi-\frac{m^{2}}{\hbar^{2}}\phi=0\leavevmode\nobreak\ ,$ (III.1)
@@
$I[x(\lambda)]=\frac{1}{2}\int_{0}^{\lambda_{0}}g_{ab}\frac{{\rm d}x^{a}}{{\rm d}\lambda}\frac{{\rm d}x^{b}}{{\rm d}\lambda}{\rm d}\lambda\leavevmode\nobreak\ .$ (III.2)
@@
$\frac{i}{\hbar^{2}}\frac{\partial\psi}{\partial\lambda}=-\Box\psi\leavevmode\nobreak\ .$ (III.3)
To see that this is equivalent to the Klein-Gordon equation (III.1), consider a solution $\phi$ to the Klein-Gordon equation. Then one can obtain a solution of the Schrodinger equation (III.3) simply by writing $\psi(x,\lambda)=e^{im^{2}\lambda}\phi(x)$. Conversely, from a solution $\psi$ to the Schrodinger equation, one can obtain a solution to the Klein-Gordon equation by Fourier-transforming, i.e.
$\phi(x)=\int_{-\infty}^{\infty}{\rm d}\lambda\leavevmode\nobreak\ e^{-im^{2}\lambda}\psi(x,\lambda)\leavevmode\nobreak\ .$ (III.4)
@@
$=\int\mathcal{D}x(\lambda)\exp\left(\frac{i}{2\hbar}\int_{0}^{\lambda_{0}}g_{ab}\frac{{\rm d}x^{a}}{{\rm d}\lambda}\frac{{\rm d}x^{b}}{{\rm d}\lambda}\right)\leavevmode\nobreak\ .$ (III.5)
@@
$\frac{i}{\hbar}\frac{\partial F}{\partial\lambda_{0}}=-\frac{1}{2}\left(\Box-\frac{R}{3}\right)F\leavevmode\nobreak\ ,$ (III.6)
@@
$\psi(x,\lambda_{0})=\int{\rm d}^{4}x^{\prime}\leavevmode\nobreak\ F(\lambda_{0},x,x^{\prime})\psi(x^{\prime},0)\leavevmode\nobreak\ .$ (III.7)
@@
$K(x,x^{\prime})=i\int_{0}^{\infty}{\rm d}\lambda_{0}\leavevmode\nobreak\ e^{-im^{2}\lambda_{0}}F(\lambda_{0},x,x^{\prime})\leavevmode\nobreak\ .$ (III.8)
@@
$\left(\Box-\frac{m^{2}}{\hbar^{2}}\right)K(x,x^{\prime})_{=}-\frac{\delta^{(4)}(x-x^{\prime})}{\sqrt{-g}}\leavevmode\nobreak\ ,$ (III.9)
@@ IV Spontaneous emission as tunneling
Particle creation by a black hole can be understood from the perspective of pair creation via quantum mechanical tunneling [35, 47, 48, 49, 50, 51, 52]. Under normal circumstances, the vacuum of a free quantum field is populated by particle-antiparticle pairs popping into existence and annihilating as soon as they are formed – the phenomenon of virtual pair-creation. Thus, on average, an observer will detect no particles. However, in the presence of a strong potentials (e.g. electromagnetic fields, gravitational fields, etc.), some of the virtual particle-antiparticle pairs might get separated from each other, and hence can eventually be registered by a particle detector – the phenomenon of real pair creation. The fact that this process cannot occur classically can be thought of in terms of there being a ``barrier'' which prevents particle-antiparticle pairs from being separated; quantum mechanically, however, there can be tunneling through a barrier, and thus the possibility suggests itself of understanding spontaneous emission by black holes as a tunneling phenomenon.
@@
$K(x,x^{\prime})=N\exp\left(\frac{i}{\hbar}I[x,x^{\prime}]\right)\leavevmode\nobreak\ ,$ (IV.1)
@@
$\partial_{a}I\partial^{a}I+m^{2}=0\leavevmode\nobreak\ .$ (IV.2)
If we keep in mind the fact that the Lagrangian $\frac{1}{2}g_{ab}\frac{{\rm d}x^{a}}{{\rm d}\lambda}\frac{{\rm d}x^{b}}{{\rm d}\lambda}$ for a massive relativistic particle of mass $m$ is equal to $m^{2}/2$, then we realize that $\partial I/\partial\lambda=-m^{2}/2$ and thus the preceding equation is nothing but the Hamilton-Jacobi equation for $I$. In other words, in (IV.1), the action is to be evaluated along paths that satisfy the Hamilton-Jacobi equation. The equation (IV.1) is the path-integral analog of the WKB approximation to the wave function in quantum mechanics [53, 54].
Now, it might appear that the presence of only classical paths in (IV.1) precludes the possibility of using this equation to study quantum mechanical tunneling. The appearance, however, is illusory. To understand this point, it is instructive to recall how tunneling through a classically forbidden region is studied in the WKB approximation to the Schrodinger equation [55, Chapter VII]. Strictly speaking, the WKB approximation breaks down near the classical turning points on the particle's trajectory. To deal with this problem, one allows the particle's path to be complex and passes around the turning points via a contour in the complex plane. This yields an approximate form of the wave function in the classically forbidden region, which is then matched with the WKB form of the wave function in the classically allowed region(s) via certain connection formulae (see Ref. [55, Chapter VII] for details). The essential point is that the analytic continuation around a classical turning point imparts an imaginary exponential contribution to the wavefunction; in this way, the amplitude-squared of the wave function gives a finite probability of transmission through the classically forbidden region. (For an application of this method to spontaneous emission by a Schwarzschild black hole, see [51, 47].)
@@
$\sigma=|K(x,x^{\prime})|^{2}\sim\exp\left(-\frac{2}{\hbar}\text{Im}[I(x,x^{\prime})]\right)\leavevmode\nobreak\ .$ (IV.3)
@@
$1-(N^{r})^{2}](\partial_{r}I)^{2}+(\partial_{t}I)^{2}+2N^{r}\partial_{r}I\partial_{t}I=0\leavevmode\nobreak\ ,$ (IV.4)
@@
$1-(N^{r})^{2}]\left(\frac{{\rm d}t}{{\rm d}\lambda}\right)^{2}+\left(\frac{{\rm d}r}{{\rm d}\lambda}\right)^{2}+2N^{r}\frac{{\rm d}t}{{\rm d}\lambda}\frac{{\rm d}r}{{\rm d}\lambda}=0\leavevmode\nobreak\ .$ (IV.5)
@@
$\frac{{\rm d}r}{{\rm d}t}$ $=\pm(1\mp N^{r})=\theta_{\pm}\leavevmode\nobreak\ ,$ (IV.6)
$p_{r}$ $=\mp\frac{p_{t}}{1\mp N^{r}}=-\frac{p_{t}}{\theta_{\pm}}\leavevmode\nobreak\ \leavevmode\nobreak\ ,$ (IV.7)
where we have used (II.4) and also replaced $\partial_{a}I$ with the momentum $p_{a}$ conjugate to $x^{a}$, i.e.,
$p_{a}\equiv\frac{\partial L}{\partial({\rm d}x^{a}/{\rm d}\lambda)}=g_{ab}\frac{{\rm d}x^{b}}{{\rm d}\lambda}=\partial_{a}I\leavevmode\nobreak\ ,$ (IV.8)
@@
$\omega\equiv-p_{a}K^{a}=-p_{t}\leavevmode\nobreak\ .$ (IV.9)
@@
$u^{a}\equiv\frac{{\rm d}x^{a}}{{\rm d}\tau}=\frac{{\rm d}x^{a}}{{\rm d}t}\frac{{\rm d}t}{{\rm d}\tau}=-\frac{K^{a}}{\theta_{+}\theta_{-}}\leavevmode\nobreak\ .$ (IV.10)
@@
$\omega_{\text{st-obs}}=-u^{a}p_{a}=-\frac{\omega}{\theta_{+}\theta_{-}}\leavevmode\nobreak\ .$ (IV.11)
@@
$p_{r}=\frac{\omega}{\theta_{\pm}}\leavevmode\nobreak\ .$ (IV.12)
@@
What should be the eventual fate of the particle and the antiparticle? The particle reaches future null infinity. The ingoing antiparticle, on the other hand, must be absorbed by the the collapsing matter at, say, $q$. Physically this can be understood as follows. Matter undergoing gravitational collapse in spherical symmetry can be thought of as a collection of infinitely many spherical shells, each labeled by a parameter $s$ and falling under the weight of all the shells inside it (see Ref. [37] for the classical and Refs. [29, 59] for the effective case). That is, the equations of motion for the shells decouple, each shell's dynamics being given by its radius $r(t,s)$ as a function of time. In a simplified model (see, e.g., Ref. [60]), we may assign an invariant proper mass $\mu$ to each shell such that the total mass of all shells in an interval $[0,s]$, where 0 labels the inner most shell, is $M(s)=\mu\int_{0}^{s}\leavevmode\nobreak\ {\rm d}s$ [^3] Here $M$ is the same mass parameter that occurs in the metric (II.1). In the $(t,s)$ coordinates, this parameter is independent of $t$ [37]. This assumption and with it the whole picture of independent shells falling under their gravity breaks down if the falling shells cross, which can happen for quite generic initial data [59]. We will speculate in the concluding section on how the physical picture we have outlined changes if shell crossings do occur.. Now, in a more realistic model of pair-creation than we have described above, the created particle/antiparticle will have a proper mass, say $m$. Since in this case,
$g_{ab}\frac{{\rm d}x^{a}}{{\rm d}\lambda}\frac{{\rm d}x^{b}}{{\rm d}\lambda}=-m^{2}\leavevmode\nobreak\ ,$ (IV.17)
@@
$I=\int\partial_{a}I{\rm d}x^{a}=\int(p_{t}{\rm d}t+p_{r}{\rm d}r)=\int(-\epsilon\omega{\rm d}t+p_{r}{\rm d}r)\leavevmode\nobreak\ ,$ (IV.18)
@@
$I=\underbrace{\int_{q}^{p}(\cdots)}_{I_{1}}+\underbrace{\int_{p}^{\infty}(\cdots)}_{I_{2}}\leavevmode\nobreak\ .$ (IV.19)
@@
$\frac{{\rm d}r}{{\rm d}t}=\theta_{+}\leavevmode\nobreak\ ,\quad p_{r}=\frac{\omega}{\theta_{+}}\leavevmode\nobreak\ .$ (IV.20)
@@
$\frac{{\rm d}r}{{\rm d}t}=\theta_{-}\leavevmode\nobreak\ ,\quad p_{r}=-\frac{\omega}{\theta_{+}}\leavevmode\nobreak\ .$ (IV.21)
@@
$I_{1}=\int_{q}^{p}\omega\left(\frac{1}{\theta_{-}}-\frac{1}{\theta_{+}}\right){\rm d}r=2\int_{q}^{p}\frac{\omega}{\theta_{+}\theta_{-}}{\rm d}r\leavevmode\nobreak\ ,$ (IV.22)
@@
$\lim_{\epsilon\to 0}\left[\int_{-\infty}^{r_{H_{I}}-\epsilon}+\Delta\int_{C_{1}}+\underbrace{\int_{r_{H_{I}}+\epsilon}^{r_{H_{I}}+\epsilon}}_{I_{1}}+\int_{C_{2}}+\int_{C_{3}}\right]=0\leavevmode\nobreak\ ,$ (IV.23)
@@
The preceding equation can be rewritten more illustratively as
@@
$=-i\pi\sum_{\Delta r_{H_{I}},r_{H_{O}}}\text{Res}\left(\frac{2\omega}{\theta_{+}\theta_{-}}\right)+\text{real terms}\leavevmode\nobreak\ ,$
the second line being an application [^4] In conventional accounts of Hawking radiation as tunneling [35], it is customary to use Feynman’s $i\epsilon$ prescription to evaluate the principal value of the action along the classically forbidden trajectory. We have avoided this route in favor of the standard treatment of principal values found in mathematical physics textbooks (e.g., Ref. [62]) on account of the latter’s greater clarity when the residues of the integrand cannot be found by mere inspection and one instead has to rely on a Laurent expansion. of the calculus of residues to poles on the real line [62]. To evaluate the residues, we can expand the integrand near the horizon along the path of the ingoing particle, i.e., along ${\rm d}r/{\rm d}t=\theta_{-}$. [35]. Writing $\delta r=r-r_{H}$, we have
@@
Here we have used (II.14) in the last line. Thus, to first order in $\delta r$,
$\frac{2\omega}{\theta_{+}\theta_{-}}\simeq-\frac{\omega_{H}}{\kappa_{H}}\frac{1}{r-r_{H}}\leavevmode\nobreak\ ,$ (IV.25)
@@
$\text{Res}\left(\frac{2\omega}{\theta_{+}\theta_{-}}\right)=-\frac{\omega_{H}}{\kappa_{H}}\leavevmode\nobreak\ ,$ (IV.26)
@@
$\text{Im}(I_{1})=\Delta\frac{\pi\omega_{H_{I}}}{\kappa_{H_{I}}}+\frac{\pi\omega_{H_{O}}}{\kappa_{H_{O}}}\leavevmode\nobreak\ .$ (IV.27)
@@
$\sigma\sim\exp\left(-\Delta\frac{2\pi\omega_{H_{I}}}{\kappa_{H_{I}}}-\frac{2\pi\omega_{H_{O}}}{\kappa_{H_{O}}}\right)\leavevmode\nobreak\ .$ (IV.28)
In the classical case of a Schwarzschild spacetime, which is characterized by an eternal event horizon, one will only have the term with $H_{O}$, which will then denote the event horizon. Furthermore, in a static spherically symmetric spacetime, the Kodama vector field reduces to the timelike Killing vector field. Therefore, $\omega=-p_{a}K^{a}$ is conserved and equals the energy $\omega_{\infty}$ of a particle as measured by an observer at infinity (since $K^{a}$ becomes a unit timelike vector field asymptotically). And finally, one can read off the surface gravity from (II.14). The preceding formula then reduces to
$\sigma\sim\exp\left(-8\pi M_{S}\leavevmode\nobreak\ \omega_{\infty}\right)\leavevmode\nobreak\ ,$ (IV.29)

### v3

@@ II.2 Dynamical horizons
By way of illustration, let us apply the formulae developed above to the framework for quantum-corrected dust collapse described in the preceding section. First of all, regardless of the explicit form of the mass function $M$, we know that it is bounded above (cf. (II.8)). Since $r_{H}=2M_{H}$, and in particular, $M_{H_{O}}=M_{\text{ADM}}$, we obtain a lower bound on the mass of the black hole:
$M_{\text{ADM}}\geq\frac{\sqrt{\mu}}{2}\penalty 10000\ .$ (II.37)
Thus, we see that a black hole forms only if the ADM mass of the collapsing body is greater than the Planck mass ($\sim\sqrt{\hbar}\sim\sqrt{\mu}$).
For the quantum-corrected Oppenheimer-Snyder model, we can calculate the surface gravity (II.34) explicitly. For the outer horizon, $\kappa_{H_{O}}=1/4M_{\text{ADM}}$, regardless of the model of collapse. Thus, it is only the surface gravity at the inner horizon $H_{I}$ which will be different for different models. For the quantum-corrected Oppenheimer-Snyder model, by setting $M_{i}(t,r)=M_{H_{I}}=r_{H_{I}}/2$ ((II.21)), we get
$M_{H_{I}}=\frac{1}{2}\sqrt{\mu+\frac{9}{4}(t-t_{0})^{2}}\penalty 10000\ ,$ (II.38)
from which we obtain
$\kappa_{H_{I}}=\frac{1}{2\sqrt{\mu+\frac{9}{4}(t-t_{0})^{2}}}\left[1+\frac{9}{4}\frac{(t-t_{0})}{\sqrt{\mu+\frac{9}{4}(t-t_{0})^{2}}}\right]\penalty 10000\ .$ (II.39)
One particular approximation to this expression will be useful later. If $(t-t_{0})\gg\mu$, which will be true soon after the bounce (since $\mu\ll 1$), we may turn off $\mu$ and write
$\kappa_{H_{I}}\approx\frac{5}{6(t-t_{0})}\penalty 10000\ .$ (II.40)
Applying the same approximation to $r_{H_{I}}=2M_{H_{I}}$, we can also write
$r_{H_{I}}=\frac{3}{2}(t-t_{0})\penalty 10000\ .$ (II.41)
One final aspect of the Oppenheimer-Snyder model that is worth bearing in mind is the following. The expression for $M_{H_{I}}$ implies that
$\frac{4\dot{M}_{H_{I}}}{(1-M^{\prime}_{H_{I}})^{2}}[1$- $(\dot{M}_{H_{I}}-2M^{\prime}_{H_{I}})]=5\dot{M}_{H_{I}}$ (II.42)
$=\frac{45}{8}\frac{t-t_{0}}{\sqrt{\mu+\frac{9}{4}(t-t_{0})^{2}}}\penalty 10000\ .$ (II.43)
It follows that during the collapse (rebound) phase, the metric on the horizon (II.29) has negative (positive) signature, confirming our expectation that the inner horizon should be timelike before the bounce and spacelike after the bounce.
II.3 Minimal assumptions on quantum-corrected collapse
It is important to emphasize that the particular framework of effective LQG and the specific example of the Oppenheimer-Snyder model were presented to illustrate concretely how the picture of quantum-corrected collapse depicted in Fig. 1 can be mathematically realized. Specific features of the idealized quantum-corrected collapse that have been presented above, such as the explicit form of the mass function and surface gravity, the speed with which the boundary of the collapsing matter moves, and so on, may turn out to be completely different if obtained from a fully consistent and satisfactory theory of quantum gravity. For the purposes of this paper, we require only the following minimal assumptions on quantum-corrected collapse in spherical symmetry, gleaned from our foregoing study of the framework of effective LQG.
  1. 1.
    Collapse and expansion occur in a single asymptotic region, as in Fig. 1. This means that the spacetime can be covered by generalized PG coordinates using the metric (II.1).
  2. 2.
    The conserved mass of the collapsing matter should be much larger than the Planck mass, $M_{\text{ADM}}\gg\sqrt{\hbar}$, consistent with (II.37). This in turn implies that the trapped region in Fig. 1 is macroscopic, i.e., $r_{H_{O}}\gg\sqrt{\hbar}$.
The justification for the first assumption is its attestation in a growing number of quantum-corrected models of gravitational collapse [67, 35, 81, 76, 10, 11, 60, 52, 53, 82, 22, 14, 8, 36, 37, 34, 33, 1, 19]. The second assumption is necessary because such models are to be understood in the framework of an effective theory which incorporates some, but by no means all, crucial aspects of quantum gravity. As such, there is no guarantee that these models give a fully reliable picture in the small and highly quantum region where the collapsing matter bounces [50]. Therefore, any analyses using such models should be performed only in regions far away from the unknown fully quantum region. This can be done only if the size of the trapping region is much larger than the quantum region where the bounce occurs.
This concludes our review of the tools required to study effective dynamical black holes. Next we review a framework in which spontaneous emission by such black holes can be conveniently described.
@@ III Quantum theory of a relativistic particle
$\Box\phi-\frac{m^{2}}{\hbar^{2}}\phi=0\penalty 10000\ ,$ (III.1)
@@
$I[x(\lambda)]=\frac{1}{2}\int_{0}^{\lambda_{0}}g_{ab}\frac{{\rm d}x^{a}}{{\rm d}\lambda}\frac{{\rm d}x^{b}}{{\rm d}\lambda}{\rm d}\lambda\penalty 10000\ .$ (III.2)
@@
$\frac{i}{\hbar}\frac{\partial\psi}{\partial\lambda}=-\Box\psi\penalty 10000\ .$ (III.3)
To see that this is equivalent to the Klein-Gordon equation (III.1), consider a solution $\phi$ to the Klein-Gordon equation. Then one can obtain a solution of the Schrodinger equation (III.3) simply by writing $\psi(x,\lambda)=e^{im^{2}\lambda/\hbar}\phi(x)$. Conversely, from a solution $\psi$ to the Schrodinger equation, one can obtain a solution to the Klein-Gordon equation by Fourier-transforming, i.e.
$\phi(x)=\int_{-\infty}^{\infty}{\rm d}\lambda\penalty 10000\ e^{-im^{2}\lambda/\hbar}\psi(x,\lambda)\penalty 10000\ .$ (III.4)
@@
$=\int\mathcal{D}x(\lambda)\exp\left(\frac{i}{2\hbar}\int_{0}^{\lambda_{0}}g_{ab}\frac{{\rm d}x^{a}}{{\rm d}\lambda}\frac{{\rm d}x^{b}}{{\rm d}\lambda}\right)\penalty 10000\ .$ (III.5)
@@
$\frac{i}{\hbar}\frac{\partial F}{\partial\lambda_{0}}=-\frac{1}{2}\left(\Box-\frac{R}{3}\right)F\penalty 10000\ ,$ (III.6)
@@
$\psi(x,\lambda_{0})=\int{\rm d}^{4}x^{\prime}\penalty 10000\ F(\lambda_{0},x,x^{\prime})\psi(x^{\prime},0)\penalty 10000\ .$ (III.7)
@@
$K(x,x^{\prime})=i\int_{0}^{\infty}{\rm d}\lambda_{0}\penalty 10000\ e^{-im^{2}\lambda_{0}/\hbar}F(\lambda_{0},x,x^{\prime})\penalty 10000\ .$ (III.8)
@@
$\left(\Box-\frac{m^{2}}{\hbar^{2}}\right)K(x,x^{\prime})_{=}-\frac{\delta^{(4)}(x-x^{\prime})}{\sqrt{-g}}\penalty 10000\ ,$ (III.9)
@@ IV Spontaneous emission as tunneling
Particle creation by a black hole can be understood from the perspective of pair creation via quantum mechanical tunneling [78, 70, 64, 18, 71, 23, 84, 13]. Under normal circumstances, the vacuum of a free quantum field is populated by particle-antiparticle pairs popping into existence and annihilating as soon as they are formed – the phenomenon of virtual pair-creation. Thus, on average, an observer will detect no particles. However, in the presence of a strong potentials (e.g. electromagnetic fields, gravitational fields, etc.), some of the virtual particle-antiparticle pairs might get separated from each other, and hence can eventually be registered by a particle detector – the phenomenon of real pair creation. The fact that this process cannot occur classically can be thought of in terms of there being a ``barrier'' which prevents particle-antiparticle pairs from being separated; quantum mechanically, however, there can be tunneling through a barrier, and thus the possibility suggests itself of understanding spontaneous emission by black holes as a tunneling phenomenon.
@@
$K(x,x^{\prime})=N\exp\left(\frac{i}{\hbar}I[x,x^{\prime}]\right)\penalty 10000\ ,$ (IV.1)
@@
$\partial_{a}I\partial^{a}I+\frac{m^{2}}{\hbar^{2}}=0\penalty 10000\ .$ (IV.2)
If we keep in mind the fact that the Lagrangian $\frac{1}{2}g_{ab}\frac{{\rm d}x^{a}}{{\rm d}\lambda}\frac{{\rm d}x^{b}}{{\rm d}\lambda}$ for a massive relativistic particle of mass $m$ is equal to $m^{2}/2$, then we realize that $\partial I/\partial\lambda=-m^{2}/2$ and thus the preceding equation is nothing but the Hamilton-Jacobi equation for II. In other words, in (IV.1), the action is to be evaluated along paths that satisfy the Hamilton-Jacobi equation. The equation (IV.1) is the path-integral analog of the WKB approximation to the wave function in quantum mechanics [61, 49].
Now, it might appear that the presence of only classical paths in (IV.1) precludes the possibility of using this equation to study quantum mechanical tunneling. The appearance, however, is illusory. To understand this point, it is instructive to recall how tunneling through a classically forbidden region is studied in the WKB approximation to the Schrodinger equation [57, Chapter VII]. Strictly speaking, the WKB approximation breaks down near the classical turning points on the particle's trajectory. To deal with this problem, one allows the particle's path to be complex and passes around the turning points via a contour in the complex plane. This yields an approximate form of the wave function in the classically forbidden region, which is then matched with the WKB form of the wave function in the classically allowed region(s) via certain connection formulae (see Ref. [57, Chapter VII] for details). The essential point is that the analytic continuation around a classical turning point imparts an imaginary exponential contribution to the wavefunction; in this way, the amplitude-squared of the wave function gives a finite probability of transmission through the classically forbidden region. (For an application of this method to spontaneous emission by a Schwarzschild black hole, see [23, 70].)
@@
$\sigma=|K(x,x^{\prime})|^{2}\sim\exp\left(-\frac{2}{\hbar}\text{Im}[I(x,x^{\prime})]\right)\penalty 10000\ .$ (IV.3)
@@
$1-(N^{r})^{2}](\partial_{r}I)^{2}+(\partial_{t}I)^{2}+2N^{r}\partial_{r}I\partial_{t}I=0\penalty 10000\ ,$ (IV.4)
@@
$1-(N^{r})^{2}]\left(\frac{{\rm d}t}{{\rm d}\lambda}\right)^{2}+\left(\frac{{\rm d}r}{{\rm d}\lambda}\right)^{2}+2N^{r}\frac{{\rm d}t}{{\rm d}\lambda}\frac{{\rm d}r}{{\rm d}\lambda}=0\penalty 10000\ .$ (IV.5)
@@
$\frac{{\rm d}r}{{\rm d}t}$ $=\pm(1\mp N^{r})=\theta_{\pm}\penalty 10000\ ,$ (IV.6)
$p_{r}$ $=\mp\frac{p_{t}}{1\mp N^{r}}=-\frac{p_{t}}{\theta_{\pm}}\penalty 10000\ \penalty 10000\ ,$ (IV.7)
where we have used (II.24) and also replaced $\partial_{a}I$ with the momentum $p_{a}$ conjugate to $x^{a}$, i.e.,
$p_{a}\equiv\frac{\partial L}{\partial({\rm d}x^{a}/{\rm d}\lambda)}=g_{ab}\frac{{\rm d}x^{b}}{{\rm d}\lambda}=\partial_{a}I\penalty 10000\ ,$ (IV.8)
@@
$\omega\equiv-p_{a}K^{a}=-p_{t}\penalty 10000\ .$ (IV.9)
@@
$u^{a}\equiv\frac{{\rm d}x^{a}}{{\rm d}\tau}=\frac{{\rm d}x^{a}}{{\rm d}t}\frac{{\rm d}t}{{\rm d}\tau}=-\frac{K^{a}}{\theta_{+}\theta_{-}}\penalty 10000\ .$ (IV.10)
@@
$\omega_{\text{st-obs}}=-u^{a}p_{a}=-\frac{\omega}{\theta_{+}\theta_{-}}\penalty 10000\ .$ (IV.11)
@@
$p_{r}=\frac{\omega}{\theta_{\pm}}\penalty 10000\ .$ (IV.12)
@@
What should be the eventual fate of the particle and the antiparticle? The particle reaches future null infinity. The ingoing antiparticle, on the other hand, must be absorbed by the the collapsing matter at, say, $q$. Physically this can be understood as follows. Matter undergoing gravitational collapse in spherical symmetry can be thought of as a collection of infinitely many spherical shells, each labeled by a parameter $s$ and falling under the weight of all the shells inside it (see Ref. [17] for the classical and Refs. [34, 26] for the effective case). That is, the equations of motion for the shells decouple, each shell's dynamics being given by its radius $r(t,s)$ as a function of time. In a simplified model (see, e.g., Ref. [31]), we may assign an invariant proper mass $\mu$ to each shell such that the total mass of all shells in an interval $[0,s]$, where 0 labels the inner most shell, is $M(s)=\mu\int_{0}^{s}\penalty 10000\ {\rm d}s$ [^8] Here $M$ is the same mass parameter that occurs in the metric (II.1). In the $(t,s)$ coordinates, this parameter is independent of $t$ [17]. This assumption and with it the whole picture of independent shells falling under their gravity breaks down if the falling shells cross, which can happen for quite generic initial data [26]. We will speculate in the concluding section on how the physical picture we have outlined changes if shell crossings do occur.. Now, in a more realistic model of pair-creation than we have described above, the created particle/antiparticle will have a proper mass, say $m$. Since in this case,
$g_{ab}\frac{{\rm d}x^{a}}{{\rm d}\lambda}\frac{{\rm d}x^{b}}{{\rm d}\lambda}=-m^{2}\penalty 10000\ ,$ (IV.17)
@@
$I=\int\partial_{a}I{\rm d}x^{a}=\int(p_{t}{\rm d}t+p_{r}{\rm d}r)=\int(-\epsilon\omega{\rm d}t+p_{r}{\rm d}r)\penalty 10000\ ,$ (IV.18)
@@
$I=\underbrace{\int_{q}^{p}(\cdots)}_{I_{1}}+\underbrace{\int_{p}^{\infty}(\cdots)}_{I_{2}}\penalty 10000\ .$ (IV.19)
@@
$\frac{{\rm d}r}{{\rm d}t}=\theta_{+}\penalty 10000\ ,\quad p_{r}=\frac{\omega}{\theta_{+}}\penalty 10000\ .$ (IV.20)
@@
$\frac{{\rm d}r}{{\rm d}t}=\theta_{-}\penalty 10000\ ,\quad p_{r}=-\frac{\omega}{\theta_{+}}\penalty 10000\ .$ (IV.21)
@@
$I_{1}=\int_{q}^{p}\omega\left(\frac{1}{\theta_{-}}-\frac{1}{\theta_{+}}\right){\rm d}r=2\int_{q}^{p}\frac{\omega}{\theta_{+}\theta_{-}}{\rm d}r\penalty 10000\ ,$ (IV.22)
@@
$\lim_{\epsilon\to 0}\left[\int_{-\infty}^{r_{H_{I}}-\epsilon}+\Delta\int_{C_{1}}+\underbrace{\int_{r_{H_{I}}+\epsilon}^{r_{H_{O}}-\epsilon}}_{I_{1}}+\int_{C_{2}}+\int_{C_{3}}\right]=0\penalty 10000\ ,$ (IV.23)
@@
The preceding equation can be rewritten more illustratively as [^9] Strictly speaking, because we are working in spherical coordinates, the integrand is not defined for $r<0$. But the term $\int_{-\infty}^{r_{H_{I}}-\epsilon}$ necessitates that we specify the integrand on the negative real axis too. To do this, we may make any choice as long as $\omega$ and $M$ are real for $r<0$. In this we are guided by the expectation that this term should not give an imaginary contribution, as is the case for a Schwarzschild black-hole, for which the integrand is $2\omega_{\infty}/(1-2M/r)$, which is real even if $r<0$. I thank Irfan Javed for drawing my attention to this subtle point.
@@
$=-i\pi\sum_{\Delta r_{H_{I}},r_{H_{O}}}\text{Res}\left(\frac{2\omega}{\theta_{+}\theta_{-}}\right)+\text{real terms}\penalty 10000\ ,$
the second line being an application [^10] In conventional accounts of Hawking radiation as tunneling [78], it is customary to use Feynman’s $i\epsilon$ prescription to evaluate the principal value of the action along the classically forbidden trajectory. We have avoided this route in favor of the standard treatment of principal values found in mathematical physics textbooks (e.g., Ref. [2]) on account of the latter’s greater clarity when the residues of the integrand cannot be found by mere inspection and instead one has to rely on a Laurent expansion. of the calculus of residues to poles on the real line [2]. To evaluate the residues, we can expand the integrand near the horizon along the path of the ingoing particle, i.e., along ${\rm d}r/{\rm d}t=\theta_{-}$. [78]. Writing $\delta r=r-r_{H}$, we have
@@
Here we have used (II.34) in the last line. Thus, to first order in $\delta r$,
$\frac{2\omega}{\theta_{+}\theta_{-}}\simeq-\frac{\omega_{H}}{\kappa_{H}}\frac{1}{r-r_{H}}\penalty 10000\ ,$ (IV.25)
@@
$\text{Res}\left(\frac{2\omega}{\theta_{+}\theta_{-}}\right)=-\frac{\omega_{H}}{\kappa_{H}}\penalty 10000\ ,$ (IV.26)
@@
$\text{Im}(I_{1})=\Delta\frac{\pi\omega_{H_{I}}}{\kappa_{H_{I}}}+\frac{\pi\omega_{H_{O}}}{\kappa_{H_{O}}}\penalty 10000\ .$ (IV.27)
@@
$\sigma\sim\exp\left(-\Delta\frac{2\pi\omega_{H_{I}}}{\hbar\kappa_{H_{I}}}-\frac{2\pi\omega_{H_{O}}}{\hbar\kappa_{H_{O}}}\right)\penalty 10000\ .$ (IV.28)
In the classical case of a Schwarzschild spacetime, which is characterized by an eternal event horizon, one will only have the term with $H_{O}$, which will then denote the event horizon. Furthermore, in a static spherically symmetric spacetime, the Kodama vector field reduces to the timelike Killing vector field. Therefore, $\omega=-p_{a}K^{a}$ is conserved and equals the energy $\omega_{\infty}$ of a particle as measured by an observer at infinity (since $K^{a}$ becomes a unit timelike vector field asymptotically). And finally, one can read off the surface gravity from (II.34). The preceding formula then reduces to
$\sigma\sim\exp\left(-8\pi M_{S}\penalty 10000\ \omega_{\infty}/\hbar\right)\penalty 10000\ ,$ (IV.29)